# Pathfinder for Low-altitude Aircraft with Binary Neural Network

### v1

@@ Abstract
A prior global topological map (e.g., the OpenStreetMap, OSM) can boost the performance of autonomous mapping by a ground mobile robot. However, the prior map is usually incomplete due to lacking labeling in partial paths. To solve this problem, this paper proposes an OSM maker using airborne sensors carried by low-altitude aircraft, where the core of the OSM maker is a novel efficient pathfinder approach based on LiDAR and camera data, i.e., a binary dual-stream road segmentation model. Specifically, a multi-scale feature extraction based on the UNet architecture is implemented for images and point clouds. To reduce the effect caused by the sparsity of point cloud, an attention-guided gated block is designed to integrate image and point-cloud features. For enhancing the efficiency of the model, we propose a binarization streamline to each model component, including a variant of vision transformer (ViT) architecture as the encoder of the image branch, and new focal and perception losses to optimize the model training. The experimental results on two datasets demonstrate that our pathfinder method achieves SOTA accuracy with high efficiency in finding paths from the low-level airborne sensors, and we can create complete OSM prior maps based on the segmented road skeletons. Code and data are available at: https://github.com/IMRL/Pathfinder.
@@ I INTRODUCTION
In recent years, autonomous mapping based on ground robots has become a popular research area [1, 2]. Existing methods can generally work well in small areas with simple topology but usually fail in large scenes [3, 4, 5, 6]. To solve this problem, some mapping methods guided by prior global maps have been proposed, such as based on the GPS coordinates of Open Street Maps (OSM)  [1, 7, 8]. With prior global information, robots can plan an optimal path to follow and complete the mapping on a large scale. However, prior maps are usually incomplete due to the loss of GPS signals or the lack of labelling in OSM (as shown in Fig.1). Thus, robots cannot scout every corner of the scene even if they follow the planned path according to the prior map.
This work addresses this issue by proposing an OSM maker using airborne sensors (LiDAR and camera). The OSM maker can generate an entire topology of traversable road networks and provide the complete prior OSM map required to guide ground robots in autonomous mapping, acting as a mapping assistant for ground robots in a cooperative mapping framework [9, 10, 11]. Like any real-time mapping and robotic exploration algorithms, the OSM maker should be as efficient as possible given the low-cost and power-efficiency requirements. Thus, we propose a novel efficient binary dual-stream road segmentation model, where weights and activations are represented with only 1 bit. For 1-bit weights and activations, Xnor and PopCount operations [12] can replace floating-point multiplication and addition (MAC) operations, significantly reducing memory requirements and computational costs. To our best knowledge, it is the first binary dual-stream model that effectively fuses RGB images and LiDAR point-cloud data for road segmentation tasks, not limited to applications to low-altitude aircraft. The contributions of this work are summarized as follows:
@@
  • We propose a novel binary dual-stream framework, Binarized Pathfinder, of accurate and efficient road segmentation for OSM generation.
@@ IV-D Ablation Study
  Network Flops (G) Bops (G) Ops (G) Para.(MB)
  PMF-Net (FP) 237.71 0 237.71 138.90
Ours (FP) 202.28 0 202.28 153.69
Ours (binary) 13.46 188.82 16.41 15.24
@@ V CONCLUSIONS
In this paper, we propose an OSM maker to tackle the problem, labelling lacking in partial paths, within the current OSM method. A key component of our approach is an efficient binary dual-stream road segmentation model built on the Unet architecture. To boost the performance of our model, we employ improved loss functions and specially designed modules for method optimization. Experimental results demonstrate that our approach attains state-of-the-art accuracy (mIOU) while maintaining low memory usage and computational complexity in finding path, which ensures the capability of OSM maker to complete or maintain high-precision OSM prior maps efficiently.

### v3

@@ Abstract
A prior global topological map (e.g., the OpenStreetMap, OSM) can boost the performance of autonomous mapping by a ground mobile robot. However, the prior map is usually incomplete due to lacking labeling in partial paths. To solve this problem, this paper proposes an OSM maker using airborne sensors carried by low-altitude aircraft, where the core of the OSM maker is a novel efficient pathfinder approach based on LiDAR and camera data, i.e., a binary dual-stream road segmentation model. Specifically, a multi-scale feature extraction based on the UNet architecture is implemented for images and point clouds. To reduce the effect caused by the sparsity of point cloud, an attention-guided gated block is designed to integrate image and point-cloud features. To optimize the model for edge deployment that significantly reduces storage footprint and computational demands, we propose a binarization streamline to each model component, including a variant of vision transformer (ViT) architecture as the encoder of the image branch, and new focal and perception losses to optimize the model training. The experimental results on two datasets demonstrate that our pathfinder method achieves SOTA accuracy with high efficiency in finding paths from the low-level airborne sensors, and we can create complete OSM prior maps based on the segmented road skeletons. Code and data are available at: https://github.com/IMRL/Pathfinder.
@@ I INTRODUCTION
In recent years, autonomous mapping based on ground robots has become a popular research area [1, 2]. Existing methods can generally work well in small areas with simple topology but usually fail in large scenes [3, 4, 5, 6]. To solve this problem, some mapping methods guided by prior global maps have been proposed, such as based on the GPS coordinates of Open Street Maps (OSM)  [1, 7, 8]. With prior global information, robots can plan an optimal path to follow and complete the mapping on a large scale. However, prior maps are usually incomplete due to the loss of GPS signals or the lack of labeling in OSM (as shown in Fig. 1). Thus, robots cannot scout every corner of the scene even if they follow the planned path according to the prior map.
This work addresses this issue by proposing an OSM maker using airborne sensors (LiDAR and camera). The integration of LiDAR and camera addresses the sparsity issue of LiDAR data while enhancing robustness to adverse lighting conditions (e.g., shadows) compared to single sensor solutions (as shown in Fig. 10). The OSM maker can generate an entire topology of traversable road networks and provide the complete prior OSM map required to guide ground robots in autonomous mapping, acting as a mapping assistant for ground robots in a cooperative mapping framework [9, 10, 11]. Like any real-time mapping and robotic exploration algorithms, the OSM maker should be as efficient as possible, given the low-cost and power-efficiency requirements. Thus, we propose an efficient binary dual-stream road segmentation model, where weights and activations are represented with only 1 bit. For 1-bit weights and activations, Xnor and PopCount operations [12] can replace floating-point multiplication and addition (MAC) operations, significantly reducing memory requirements and computational costs. To our best knowledge, it is the first binary dual-stream model that effectively fuses RGB images and LiDAR point-cloud data for road segmentation tasks, not limited to applications to low-altitude aircraft. The contributions of this work are summarized as follows:
@@
  • We propose a novel binary dual-stream framework, Binarized Pathfinder, of accurate and efficient road segmentation for OSM generation. Additionally, we propose the RS-LVF dataset to support this task.
@@ IV-D Ablation Study
  Network Ops (G) Para. (MB) Latency (ms)
  PMF-Net (FP) 237.71 138.90 1787
Ours (FP) 202.28 153.69 1754
Ours (binary) 16.41 15.24 604
@@
Latency result on edge devices. To measure real-time hardware latency, we first convert the PyTorch implementation of the proposed framework to ONNX format and then deploy it on an ARM Cortex-A76-based edge device (without CUDA) using the Bolt toolkit. Experimental results in Table IV demonstrate that the 1-bit version reduces latency by 2.9× compared to full-precision methods. We anticipate further acceleration potential through dedicated hardware accelerators optimized for binarized operations.
@@ V CONCLUSIONS
In this paper, we propose an OSM maker to address missing partial path labels in existing OSM methods. The core innovation is a UNet-based binary dual-stream road segmentation model, enhanced with optimized loss functions and specialized modules. Experiments demonstrate state-of-the-art mIOU with low memory/computational costs, enabling efficient high-precision OSM map completion and maintenance.